OpenStudy (anonymous):

In how many ways can a teacher arrange five students in the front row of a classroom with a total of 20 students? Answer 860,480 2,860,480 60,480 1,860,480

OpenStudy (anonymous):

\[20\times 19\times 18\times 16\times 15\]

OpenStudy (tkhunny):

17?